OpenStudy (chris215):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y=2/(x^2)+3

OpenStudy (anonymous):

one way is that you take g(x)=x^2 and f(x)=2/x + 3. then y=f(g(x))= f(x^2)= y=2/x^2 + 3

OpenStudy (anonymous):

it's a very easy technique. just take g(x) to be anything relating to the expression in the equation.

OpenStudy (anonymous):

for example, in this equation, you can take g(x) to be x^2, 1/x, 1/x^2, etc.

OpenStudy (chris215):

I'm really confused could you explain it a little more?

OpenStudy (anonymous):

yeah of course

OpenStudy (anonymous):

Now, you have choose f(x) accordingly. if you take g(x)=x^2, then you have to take f(x)=2/x + 3; if you take g(x)=1/x, then you have to take f(x)=2x^2+3, and so on.

OpenStudy (anonymous):

you may ask your self, but is that the answer to what f(x) and g(x) are??

OpenStudy (anonymous):

there are infinite no. of functions that can work as f and g. there is no one answer.

OpenStudy (anonymous):

for example, you take g(x)=2/x^(2) and f(x)=x+3; you can take g(x)=2/x^(2) + 1 and f(x)=x+2, g(x)=2/x^(2)+2 and f(x)=x+1... these all satisfy y=f(g(x)).

OpenStudy (anonymous):

did i help?

OpenStudy (anonymous):

cause thats pretty much all ik

OpenStudy (chris215):

ohhh ok I think I understand a bit better thanks so much

OpenStudy (anonymous):

np glad i could help you :)